# BIS Reasoning 1.0: The First Large-Scale Japanese Benchmark for Belief-Inconsistent Syllogistic Reasoning

### v4

@@ 4.4 Detailed Error Analysis for GPT-4o
Conversely, informal prompts (casual) and simple instruction (basic) achieved extremely low recovery rates (3% and 5%, respectively). These prompts appear insufficient in addressing GPT-4o’s belief bias, indicating that the model defaults to common-sense heuristics without clear guidance.
The polite ({CJK}UTF8ipxm敬語, keigo) prompt yielded a moderate accuracy improvement (65%), suggesting linguistic politeness cues may moderately encourage GPT-4o to engage in deeper reflection or careful reasoning, though less effectively than explicitly logical or CoT framing.
Repeating this experiment with English-language prompts while keeping the Japanese data constant yielded a similar performance pattern, as illustrated in Figure 4. However, the accuracy gaps were less pronounced compared to the Japanese prompts. This narrower gap in performance likely stems from GPT-4o being extensively trained in English, enhancing its robustness to varied prompt styles. Nevertheless, this confirms the key insight: prompt design significantly influences GPT-4o’s performance on challenging tasks like belief-inconsistent syllogisms.
Figure 4: Prompt Type Accuracy on Error Samples (%) – Retest with English prompts
Table 2 provides detailed descriptions of each prompt type used in the error re-evaluation.
Table 2: Descriptions of prompt types used for re-evaluating GPT-4o errors.
@@
These findings demonstrate the substantial impact of prompt engineering on GPT-4o’s logical reasoning capabilities, particularly in overcoming belief bias. Although GPT-4o already performs strongly relative to other models, explicit prompting—such as instructing it to follow a structured reasoning process or clearly signaling the presence of belief-inconsistent content—markedly enhances its logical consistency. Consequently, when deploying LLMs, especially in contexts demanding precise and unbiased logical inference, strategic prompt design is crucial. Clear, structured instructions, emphasizing logical rigor and explicitly guiding step-by-step reasoning, can significantly mitigate intuitive biases inherent to the model, thus ensuring more reliable and accurate outcomes.
4.5 Logit Lens Analysis for Open Models
Based on the results in Table 1, we observe that the open model llm-jp-3-13b achieves competitive performance compared to commercial alternatives. To gain deeper insights into the model’s internal state during the belief-inconsistent syllogistic reasoning task, we conduct an analysis using the logit lens technique.
We perform this analysis on two models: llm-jp/llm-jp-3-13b and stockmark/stockmark-13b. Both are open, Japanese-optimized models of equal size (13B parameters), making them ideal for controlled comparison. Figure 5 presents a representative example of the logit lens output.
The key observation is the model’s behavior at the Conclusion_End token position. For stockmark/stockmark-13b, its prediction probability becomes high early in the network, suggesting a form of premature confidence. In contrast, llm-jp/llm-jp-3-13b shows a gradual increase in this signal across layers, indicating a delayed and layered buildup of reasoning. Additionally, the Premise1_End signal is notably weaker in stockmark, implying possible uncertainty in boundary detection between the premises and the conclusion.
(a) stockmark/stockmark-13b
(b) llm-jp/llm-jp-3-13b
Figure 5: Logit lens visualization on the belief-inconsistent syllogistic reasoning task.
5 Discussion
The results of our evaluation highlight important distinctions in the reasoning abilities of contemporary language models when confronted with belief-inconsistent syllogisms. GPT-4o’s leading performance—nearly 80% accuracy—suggests that advanced training strategies, possibly including large-scale instruction tuning and reasoning-augmented supervision, have enabled it to prioritize formal logic over commonsense intuition. Its effectiveness on a Japanese-language task, despite not being explicitly tailored for Japanese, underscores the strength of cross-lingual transfer when reasoning is properly grounded.
In contrast, the poor performance of models like Claude-3-opus raises concerns about reasoning suppression in alignment-optimized architectures. While such models may be calibrated to avoid generating implausible outputs, this safety-oriented behavior appears to interfere with their ability to recognize and validate logically correct but counterintuitive conclusions. Similarly, the moderate but consistent performance of Japanese-specialized models suggests that strong language modeling does not necessarily imply strong reasoning. These models handle Japanese fluently yet often fail to maintain deductive consistency, particularly when logic conflicts with expectation.
Another notable finding is the absence of a positive correlation between model size and reasoning performance. Larger models in our evaluation, such as llm-jp-3-172b, underperformed their smaller counterparts, indicating that scale alone is not sufficient. Rather, performance on belief-inconsistent reasoning tasks appears to depend more on the nature of pretraining data, task-aligned fine-tuning, and perhaps architectural inductive biases that support structured inference.
Overall, our results reinforce the view that logical reasoning remains a distinct capability in language models—one that must be explicitly tested and optimized for, particularly in high-stakes domains where trust in correctness cannot rest solely on fluency or apparent plausibility.
6 Limitations
@@
In terms of model coverage, while we include both general-purpose and Japanese-specialized LLMs, our selection remains limited. Notably, open-source models outside of the LLM-JP and Stockmark ecosystems, as well as mid-sized multilingual models, are not represented in this evaluation.
@@
Lastly, model capabilities are evolving rapidly. The results presented reflect the state of model behavior at a specific point in time (early 2025), and future model updates could yield different performance profiles.
7 Conclusion
We introduced BIS Reasoning 1.0, the first large-scale Japanese-language dataset explicitly focused on belief-inconsistent syllogistic reasoning. Through a comprehensive evaluation of state-of-the-art language models, we demonstrated that reasoning under belief conflict remains a fundamental weakness—even for models with native-level Japanese capabilities. Among the models tested, GPT-4o achieved the highest accuracy, showing that architectural and training sophistication can help mitigate (though not eliminate) this form of reasoning bias.
Our results underscore a critical insight: language fluency does not guarantee logical soundness. The performance gap between model families challenges the assumption that scale or localization alone ensures robustness, and reinforces the importance of domain-targeted benchmarks for high-stakes applications. BIS Reasoning 1.0 fills this gap by offering a linguistically natural, cognitively demanding benchmark designed to expose belief-driven reasoning failures.
This work lays the groundwork for future research on belief bias mitigation, logic-aware model design, and dataset extensions across new reasoning paradigms and specialized domains. As LLMs become more deeply embedded in fields such as law, healthcare, and policy, BIS Reasoning 1.0 provides a practical tool for stress-testing their logical fidelity in scenarios where correctness must outweigh intuition.
Acknowledgements
We would like to thank Yusuke Oda, Kiyomaru Hirokazu, Maki Matsuda, and Kouta Nakayama, among others, for their valuable input—ranging from ideas for evaluating large language models, suggestions that improved the literature review, to support in drafting annotation guidelines.

### v5

@@ 4.3.  Detailed Error Analysis for GPT-4o
Conversely, informal prompts (casual) and simple instruction (basic) achieved extremely low recovery rates (3% and 5%, respectively). These prompts appear insufficient in addressing GPT-4o’s belief bias, indicating that the model defaults to commonsense heuristics without clear guidance.
The polite prompt yielded a moderate accuracy improvement (65%), suggesting linguistic politeness cues may moderately encourage GPT-4o to engage in deeper reflection or careful reasoning, though less effectively than explicitly logical or CoT framing.
Repeating this experiment with English-language prompts while keeping the Japanese data constant yielded a similar performance pattern, as illustrated in Fig.˜5. However, the accuracy gaps were less pronounced compared to the Japanese prompts. This narrower gap in performance likely stems from GPT-4o being extensively trained in English, enhancing its robustness to varied prompt styles. Nevertheless, this confirms the key insight: prompt design significantly influences GPT-4o’s performance on challenging tasks like belief-inconsistent syllogisms.
Table˜2 provides detailed descriptions of each prompt type used in the error re-evaluation.
@@
These findings demonstrate the substantial impact of prompt engineering on GPT-4o’s logical reasoning capabilities, particularly in overcoming belief bias. Although GPT-4o already performs strongly relative to other models, explicit prompting – such as instructing it to follow a structured reasoning process or clearly signaling the presence of belief-inconsistent content – markedly enhances its logical consistency. Consequently, when deploying LLMs, especially in contexts demanding precise and unbiased logical inference, strategic prompt design is crucial. Clear, structured instructions, emphasizing logical rigor and explicitly guiding step-by-step reasoning, can significantly mitigate intuitive biases inherent to the model, thus ensuring more reliable and accurate outcomes.
5.  Discussion
Our findings reveal persistent weaknesses in belief-inconsistent reasoning across LLMs. Models optimized for reasoning, such as Qwen3-32B and GPT-5, maintain high logical fidelity, while Japanese-specialized or alignment-heavy models often favor believable but invalid conclusions. This confirms that linguistic fluency and reasoning robustness are distinct capabilities. However, the latest llm-jp-3.1-13b-instruct4 marks a clear improvement over earlier llm-jp variants, indicating that recent Japanese models are beginning to incorporate stronger reasoning objectives. This trend aligns with the global movement toward reasoning-centric fine-tuning observed in post-2024 LLM generations.
Prompt design and inference-time reasoning effort strongly affect outcomes. Explicit logical instructions or chain-of-thought prompting significantly reduce belief bias, whereas minimal reasoning settings lead to sharp accuracy drops. Alignment-focused training, as seen in some Claude models, may further suppress acceptance of counterintuitive yet valid conclusions.
Model scale alone does not ensure logical reliability. Instead, reasoning-oriented training and architecture play a decisive role. These insights emphasize the need for benchmarks like BIS Reasoning 1.0 that expose belief bias and test logic over intuition. For reliable deployment in law, healthcare, and research, LLMs must be evaluated and optimized for strict logical consistency, not just fluency or alignment.
6.  Limitations
@@
In terms of model coverage, while we include both general-purpose and Japanese-specialized LLMs, our selection remains limited. Notably, open-weights models outside of the LLM-jp and Stockmark ecosystems, as well as mid-sized multilingual models, are not represented in this evaluation.
@@
Lastly, model capabilities are evolving rapidly. The results presented reflect the state of model behavior at a specific point in time (early-to-mid 2025), and future model updates could yield different performance profiles.
7.  Conclusion
We introduced BIS Reasoning 1.0, the first large-scale Japanese benchmark explicitly designed to evaluate belief-inconsistent syllogistic reasoning in large language models. Our experiments show that even advanced systems still struggle when logical validity conflicts with intuitive or factual beliefs, indicating that belief bias remains a fundamental limitation across architectures and training paradigms.
Among all evaluated models, reasoning-optimized systems such as Qwen3-32B and GPT-5 achieved near-human logical consistency, while earlier Japanese-specialized models performed significantly lower despite superior linguistic fluency. The marked improvement of llm-jp-3.1-13b-instruct4 suggests that recent Japanese LLMs are beginning to integrate reasoning-aligned fine-tuning, echoing global trends in reasoning-centric model design.
BIS Reasoning 1.0 fills a critical gap in Japanese-language reasoning evaluation by providing a natural, belief-challenging benchmark that exposes logical bias and tests reasoning fidelity under cognitive conflict. We hope this work accelerates the development of bias-resistant, logic-grounded LLMs and fosters research into reasoning alignment for safety-critical applications in law, medicine, and science.
8.  Bibliographical References
  • A. Aizawa, E. Aramaki, B. Chen, F. Cheng, H. Deguchi, R. Enomoto, K. Fujii, K. Fukumoto, T. Fukushima, N. Han, et al. (2024) Llm-jp: a cross-organizational project for the research and development of fully open japanese llms. arXiv preprint arXiv:2407.03963. Cited by: §4.1.
  • R. Ando, T. Morishita, H. Abe, K. Mineshima, and M. Okada (2023) Evaluating large language models with NeuBAROCO: syllogistic reasoning ability and human-like biases. In Proceedings of the 4th Natural Logic Meets Machine Learning Workshop, S. Chatzikyriakidis and V. de Paiva (Eds.), Nancy, France, pp. 1–11. External Links: Link Cited by: §2.
  • Y. Bai, A. Jones, K. Ndousse, A. Askell, A. Chen, N. DasSarma, D. Drain, S. Fort, D. Ganguli, T. Henighan, et al. (2022) Training a helpful and harmless assistant with reinforcement learning from human feedback. arXiv preprint arXiv:2204.05862. Cited by: §2.
  • P. Bajpai, N. Chatterjee, S. Dutta, and T. Chakraborty (2024) Can llms replace neil degrasse tyson? evaluating the reliability of llms as science communicators. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, pp. 15895–15912. Cited by: §1.
  • I. Dasgupta, A. K. Lampinen, S. C. Chan, H. R. Sheahan, A. Creswell, D. Kumaran, J. L. McClelland, and F. Hill (2022) Language models show human-like content effects on reasoning tasks. arXiv preprint arXiv:2207.07051. Cited by: §1.
  • J. S. B. Evans, J. L. Barston, and P. Pollard (1983) On the conflict between logic and belief in syllogistic reasoning. Memory & cognition 11 (3), pp. 295–306. Cited by: §2.
  • J. Frohberg and F. Binder (2022) CRASS: a novel data set and benchmark to test counterfactual reasoning of large language models. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, N. Calzolari, F. Béchet, P. Blache, K. Choukri, C. Cieri, T. Declerck, S. Goggi, H. Isahara, B. Maegaard, J. Mariani, H. Mazo, J. Odijk, and S. Piperidis (Eds.), Marseille, France, pp. 2126–2140. External Links: Link Cited by: §1.
  • T. Kojima, S. S. Gu, M. Reid, Y. Matsuo, and Y. Iwasawa (2022) Large language models are zero-shot reasoners. Advances in neural information processing systems 35, pp. 22199–22213. Cited by: §2.
  • J. Li, L. Yu, and A. Ettinger (2023) Counterfactual reasoning: testing language models’ understanding of hypothetical scenarios. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), A. Rogers, J. Boyd-Graber, and N. Okazaki (Eds.), Toronto, Canada, pp. 804–815. External Links: Link, Document Cited by: §1.
  • H. Liu, J. Liu, L. Cui, Z. Teng, N. Duan, M. Zhou, and Y. Zhang (2023) LogiQA 2.0-an improved dataset for logical reasoning in natural language understanding. IEEE ACM Trans. Audio Speech Lang. Process.. Cited by: §2.
  • T. Morishita, A. Yamaguchi, G. Morio, H. Tomonari, O. Imaichi, and Y. Sogawa (2024) JFLD: a Japanese benchmark for deductive reasoning based on formal logic. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), N. Calzolari, M. Kan, V. Hoste, A. Lenci, S. Sakti, and N. Xue (Eds.), Torino, Italia, pp. 9526–9535. External Links: Link Cited by: §1, §1, §2.
  • H. T. Nguyen, R. Goebel, F. Toni, K. Stathis, and K. Satoh (2023) A negation detection assessment of gpts: analysis with the xnot360 dataset. arXiv preprint arXiv:2306.16638. Cited by: §1.
  • K. Ozeki, R. Ando, T. Morishita, H. Abe, K. Mineshima, and M. Okada (2024) Exploring reasoning biases in large language models through syllogism: insights from the neubaroco dataset. In Findings of the Association for Computational Linguistics ACL 2024, pp. 16063–16077. Cited by: §1, §2, §2.
  • L. Qin, A. Bosselut, A. Holtzman, C. Bhagavatula, E. Clark, and Y. Choi (2019) Counterfactual story reasoning and generation. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), K. Inui, J. Jiang, V. Ng, and X. Wan (Eds.), Hong Kong, China, pp. 5043–5053. External Links: Link, Document Cited by: §1.
  • T. Sugimoto, Y. Onoe, and H. Yanaka (2023) Jamp: controlled Japanese temporal inference dataset for evaluating generalization capacity of language models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 4: Student Research Workshop), V. Padmakumar, G. Vallejo, and Y. Fu (Eds.), Toronto, Canada, pp. 57–68. External Links: Link, Document Cited by: §2.
  • J. Wei, X. Wang, D. Schuurmans, M. Bosma, F. Xia, E. Chi, Q. V. Le, D. Zhou, et al. (2022) Chain-of-thought prompting elicits reasoning in large language models. Advances in neural information processing systems 35, pp. 24824–24837. Cited by: §2.
  • C. Yan, X. Fu, Y. Xiong, T. Wang, S. C. Hui, J. Wu, and X. Liu (2025) LLM sensitivity evaluation framework for clinical diagnosis. In Proceedings of the 31st International Conference on Computational Linguistics, O. Rambow, L. Wanner, M. Apidianaki, H. Al-Khalifa, B. D. Eugenio, and S. Schockaert (Eds.), Abu Dhabi, UAE, pp. 3083–3094. External Links: Link Cited by: §1.
  • H. Yanaka and K. Mineshima (2021) Assessing the generalization capacity of pre-trained language models through japanese adversarial natural language inference. In Proceedings of the Fourth BlackboxNLP Workshop on Analyzing and Interpreting Neural Networks for NLP, pp. 337–349. Cited by: §2.
  • W. Yu, Z. Jiang, Y. Dong, and J. Feng (2025) ReClor: a reading comprehension dataset requiring logical reasoning. In International Conference on Learning Representations, Cited by: §2.